# ref.bib **footnotetext: Equal contribution

### v1

@@ D.3 Evaluating Video-LMMs
System Prompt: You are a helpful and intelligent multi-modal AI assistant, capable of performing visual question answering (VQA) task. You will be given as input 10 consecutive frames from a video, and a corresponding question related to the video, you have to answer the given question after analyzing and understanding the given input video. The question itself will present you with 4 lettered options like A) B) C) D), your task is to only output single letter corresponding to the correct answer (i.e. string literal ’A’, ’B’, ’C’, or ’D’), and you should not output anything else.
@@ Appendix E Distribution of VANE-Bench dataset
How to view the dataset? Currently, the dataset alongside metadata is hosted on the Hugging Face platform for download at this link: https://huggingface.co/datasets/rohit901/VANE-Bench. Users can directly load the dataset using Hugging Face Datasets library or download the zip file in the same Hugging Face repository. All instructions and code files to reproduce the experiments of the paper are provided in the github repository: https://github.com/rohit901/VANE-Bench/tree/main
@@
Dataset License. This work and dataset is licensed under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License. The videos in VANE-Bench dataset are collected from publicly available sources and existing real-world datasets and are for academic research use only. By using VANE-Bench, you agree not to use the dataset for any harm or unfair discrimination. Please note that the data in this dataset may be subject to other agreements. Video copyrights belong to the original dataset providers, video creators, or platforms.
URL to Croissant metadata record documenting the dataset/benchmark available for viewing and downloading by the reviewers: https://huggingface.co/api/datasets/rohit901/VANE-Bench/croissant
Appendix F Limitations
Our VANE-Bench is the first benchmark for evaluating Video-LMMs on anomalous videos from both AI-generated and real-world sources. While we have done our best to ensure a high quality evaluation of these Video-LMMs, certain limitations still manifest.
Our Question-answer pairs are designed to have 4 options. We design the instruct prompt to ensure that each Video-LMM outputs one out of 4 options. However, in some instances, model outputs hallucinated response and do not follow the instructions. As a result, we employ a post-response human-based filtration process, which involves an exhaustive verification and rectification of these errors. In our current setup, we mark these cases as wrong. We believe that future Video-LLMs will be more aligned with human intent and will follow the human instructions appropriately.
Additionally, the video samples from the SORA are limited in VANE-Bench. This is due to the fact that SORA model is not open-source yet, hence we really on publicly available samples of SORA for evaluation.

### v2

@@ D.3 Evaluating Video-LMMs
System Prompt: You are a helpful and intelligent multi-modal AI assistant, capable of performing visual question-answering (VQA) tasks. You will be given as input 10 consecutive frames from a video, and a corresponding question related to the video, you have to answer the given question after analyzing and understanding the given input video. The question itself will present you with 4 lettered options like A) B) C) D), your task is to only output single letter corresponding to the correct answer (i.e. string literal ’A’, ’B’, ’C’, or ’D’), and you should not output anything else.
@@ Appendix E Distribution of VANE-Bench dataset
How to view the dataset? The dataset alongside metadata will be hosted on the Hugging Face platform for download post acceptance of the paper. Users can directly load the dataset using Hugging Face Datasets library or download the zip file in the same Hugging Face repository. All instructions and code files to reproduce the experiments of the paper will be provided in a github repository.
@@
Dataset License. This work and dataset is licensed under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License. The videos in the VANE-Bench dataset are collected from publicly available sources and existing real-world datasets and are for academic research use only. The video generative models used to synthesize data samples in our VANE-Bench benchmark are open to use publicly and do not pose any privacy concerns as the persons or objects present in the generated videos are synthetic and do not exist in the real world. The real-world surveillance datasets - UCFCrime Sultani et al. (2018), UCSD Pedestrian Li et al. (2014), Avenue Lu et al. (2013); on the other hand, used in our work are all existing well-known and publicly available datasets that are released under open-source licenses. Thus, the original creators of these datasets have collected the data after taking informed consent from the stakeholders. By using VANE-Bench, you agree not to use the dataset for any harm or unfair discrimination. Please note that the data in this dataset may be subject to other agreements. Video copyrights belong to the original dataset providers, video creators, or platforms.